# Natal, Brazil

### v1

@@ 16 Vulnerability scan of the resulting docker image 1
Our findings corroborates the results from RQ1. In RQ1, we found that $237285\frac${237}{285}divide start_ARG 237 end_ARG start_ARG 285 end_ARG (or 83%) of the monitoring opportunities suggested the monitoring of “Coverage,” of which $205237\frac${205}{237}divide start_ARG 205 end_ARG start_ARG 237 end_ARG (or 86%) were associated with the CodeCov tool. Sonarcloud was the third most used tool among survey respondents, which is probably why “Code Quality” and “Security” practices represent the second and third strongest evidence of monitoring in PR comments, with respectively 25 and 17 mentions. Code quality and security are typically monitored by tools such as Sonarcloud.
Evaluation of CI practices
@@
When asked about the importance of monitoring the CI practices presented in our survey (see Figure 2), most participants $2028\frac${20}{28}divide start_ARG 20 end_ARG start_ARG 28 end_ARG (or 71%), provided favourable responses. Developer 7 stated: “Yes, these metrics will provide valuable information in the status of our project over time.” Developer 15: “yes, it can make the development more efficient with high quality”, Developer 24: “Yes, especially build duration, build health, and time to fix broken builds”, Developer 20: “Absolutely, it would help show the health of the project, both for the active maintainers, but also for new people evaluating the project to determine if they want to use it..”
@@
Other participants found the practices useful but not critical for their projects. Developer 9 mentioned: “In other projects these measurements are important, but in this Open Source project most of the contributors just work on the project when they have some spare time. The information is for sure useful, but not critical for us.” Some participants found only part of the practices useful. Developer 25 mentioned: “depends on the practice, for example ‘commit frequency’ is not important at all, but fixing a broken build is.” Only 2 developers found the information about CI practices evolution not useful at all. Developer 26 expressed: “Not really. I don’t think they are important because I know what needs to get done and time is the biggest limit.”, while Developer 21 said an emphatic “No.”
@@
We presume that Developer 21 is a relatively inexperienced developer. He stated that he has 05-10 years of experience with software development but is not involved in the maintenance of the CI infrastructure of the project, and he does not know which tools are involved in the project’s CI. He classified his project as having a “Base” level of CI maturity. However, because he was not involved in the project’s CI configurations and is not familiar with the CI tools used, his answers may not be entirely reliable.
@@
Developer 26 has 10-15 years of experience with software development and classified his project as having an “Intermediate” CI maturity level. His project is an embedded JavaScript library designed to include a special kind of graphics in html pages and provide them with extra functionality. Due to the nature of the project (e.g., it not a high-scale critical project), following all the CI practices may not be deemed necessary, what justifies his answer.
@@
Almost half of respondents $124\frac{11}${24}divide start_ARG 11 end_ARG start_ARG 24 end_ARG (or 46% ) would like to improve their CI practices. Developer 8 stated: “Absolutely! It saves time in the long run and increases confidence when refactoring is applied.”, Developer 28: “Yes, especially build time, fast CI = fast tests and deploy”, Developer 19: “Yes. Noticing problems early generally requires less engineering time and creates less distractions compared to noticing quality problems later.”, Developer 11: “Build time is something we should definitely worth dedicating time to. Coverage seems okay for us. Comments per change is not that much of an important aspect for us. We require at least two reviewers to approve and usually at our scale, we are positive that the reviewers usually do a good job.”, Developer 2: “Yes. Tracking the evolution of practices over time helps a project identify areas for improvement, evaluate changes, ensure consistency, and demonstrate progress to stakeholders.” and Developer 4: “Yes, as it affects the quality of our software product.”
@@
While some participants believe that it is important to invest time in improving the CI practices presented in our survey, others $524\frac${5}{24}divide start_ARG 5 end_ARG start_ARG 24 end_ARG ( or 21% ) do not consider it a priority. For example, Developer 22 stated: “Sure, if someone has time”, while Developer 9 explained: “In this open source project, it is not critical because we have people who tend to fix issues promptly. However, in teams creating software for a company, it is much more important to keep the team productive and avoid being hindered by failures in CI.”
@@
Some participants $524\frac${5}{24}divide start_ARG 5 end_ARG start_ARG 24 end_ARG ( or 21% ) do not believe that improving these CI practices are important. Developer 16 explained: “For this project, it doesn’t make much sense because the current approach seems to be working well”, while Developer 26 stated: “I don’t see how these metrics are connected to important goals like reducing errors”
@@
When asked if they find regular updates on the CI practices useful, 68% of the developers $1928\frac${19}{28}divide start_ARG 19 end_ARG start_ARG 28 end_ARG answered “Yes”, as shown in Figure 11. The responses indicate that the majority of developers consider the studied CI practices important and would monitor them if provided with the opportunity.
Figure 11: Are regular updates useful?
@@
In Figure 12, we code the benefits listed by developers when we asked about the importance of monitoring the CI practices presented in our survey. They were grouped in three main themes: (i) “benefits for the project”, (ii) “benefits for the product” and (iii) “benefits for the process”.
Figure 12: Benefits of monitoring CI practices
@@
Our findings related to this theme also corroborates our results from RQ1, where we primarily identified opportunities to monitor “build health” and “time to fix a broken build.” These two practices were considered important by developers. For example, Developer 24 stated: “Yes, especially build duration, build health, and time to fix broken builds, while Developer 20 mentioned: “Absolutely. It would help show the health of the project, both for the active maintainers and also for new people evaluating the project to determine if they want to use it.”
@@
Although most developers feel that CI impacts productivity and quality of projects, few developers $218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) found the statistical correlations between CI practices and bug-related issues or number of closed PRs useful (see Figure 3). Developer 16: “Hard to say but it makes sense that more commits lead to more closed pull requests” and Developer 13: “Generally one commit = one pull request, so I don’t think this particular metric means anything for this project.” Developer 24: “No, I did not find these 3 correlations interesting or meaningful. Not sure it will help somehow.”
@@
Only 2 developers perceived advantages in the correlations between CI practices and quality attributes presented in the survey: Developer 28: “Yes, diagnostic reasons” and Developer 20: “Yes, it shows that the CI pipeline is helpful.”
@@
In this last part of the survey, we ask participants whether there were any CI practices we missed or whether our participants preferred other types of visualizations for the metrics presented.
@@
$218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) developers expressed that the practices and visualizations presented were adequate. Developer 2: ‘Line charts can show trends in practice performance over time, while bar charts can compare the performance of different practices. Both are useful for a report on CI practices.” and Developer 27: “graphs were perfect.” $318\frac${3}{18}divide start_ARG 3 end_ARG start_ARG 18 end_ARG (or 16% ) did not miss any information, they answered a simple: “No.” $218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) developers are not sure which practices they would like to see monitored. Developer 10: “Not sure” and Developer 22: “I’m not sure.” $118\frac${1}{18}divide start_ARG 1 end_ARG start_ARG 18 end_ARG (or 5% ) Developer provided an unhelpful answer and has been discarded: “CI environments are sometimes really annoying to debug.” $1018\frac${10}{18}divide start_ARG 10 end_ARG start_ARG 18 end_ARG (or 55% ) suggested improvements to the information contained in the survey. We describe some of the suggestions in the next few paragraphs.
@@
Developer 15, while liked the visualization of the practices in the survey: “The pictures above including these 7 charts are interesting to me.”, suggested including security practices like image scan, code dependency scan, and configuration scan. Although important, these security practices are outside the scope of our survey as they do not directly relate to CI but rather to DevOps. Developer 28 suggested including “deploy time” as an interesting practice for CD in future research.
@@
Participants also discussed other matters related to PR, Issue, and Commit data. Developer 11 stated, “PR and issue related data seems to be missing. It should be public, so you should be able to gather that from our GitHub project. We use tags to sort them.” Developer 13 suggested, “Perhaps commit size over time too? Tracking keywords used in commit messages?.” Developer 12 also like how the information of seven CI practices were presented in the survey: “I think the above information is enough. For our current usage, having these simple graphs available for a daily/weekly tracking of the project status would be sufficient to assign someone to fixing the problem. We do this in a more informal way currently”, but recommended investigating “statistics on who reviews and who commits, if there’s a change in the balance (less committers or less reviewers) as important. The aim is to ensure knowledge sharing, and that one developer does not ’take over the project’ while breaking the information and knowledge sharing that is in place.” While PR and Issue information were used to compute correlations regarding CI practices in our study, they are not CI practices and were not the focus of this paper. Both “Commit size” and “Ratio between committers and reviewers,” as well the consequences of an unbalanced ratio, are interesting aspects to investigate in future research.
@@
For a broader view of the monitoring needs, in Figure 13 we show the themes that emerged from the monitoring needs suggested by survey participants.
Figure 13: Themes that emerged from Monitoring Needs
@@
Figure 14 shows the themes that emerge from the official documentation of the seven most popular CI services GolzadehCIServer2022 . The number in the center of each node indicates the number of occurrences of each theme.
@@
In this analysis we find other practices mentioned outside the scope of CI, they were classified under the umbrella of “DevOps practices”. For example, the DORA practices DoraMetricsLeanIX2023 have been widely mentioned for the CI services. Gitlab, GHActions, and Circle CI cite these practices. However, only the Gitlab documentation declare that there is an implementation of a dashboard to monitor these practices. Recently, in May 22, 2023111111https://about.gitlab.com/releases/2023/05/22/gitlab-16-0-released/, Gitlab has added the “Value Stream Daskborad” which included the DORA and Vulnerabilities metrics, but is still accessible only from Premium version [^12] https://stackoverflow.com/questions/76382478/permission-denied-for-gitlab-dora-metrics-through-api. Gitlab was the service with the highest number of citations ($49162\frac${49}{162}divide start_ARG 49 end_ARG start_ARG 162 end_ARG) (or 30% ) for themes that refer to DevOps metrics monitoring, 1.53 times more than the second place the Cicle CI with ($32162\frac${32}{162}divide start_ARG 32 end_ARG start_ARG 162 end_ARG) (or 19% ). This indicates that Gitlab is the most advanced CI service in terms of supporting the monitoring of DevOps practices, without the need to use third-party tools.
In the GHActions documentation, we did not find indications of native monitoring tools specifically designed for CI metrics. The existing monitoring tools appear to concentrate on optimizing the CI/CD pipeline by monitoring aspects such as CPU and memory usage within the pipeline. However, GHActions allows the creation of various plugins [^13] https://github.com/marketplace?category=continuous-integration&query=&type=&verification= that can complement these monitoring tasks.
Circle CI focuses on monitoring four practices: 1) “Build Health”, 2) “Mean Time to Recovery”, 3) “Build Duration”, and 4) “Build Activity”. In January 2022, a centralized dashboard [^14] https://circleci.com/docs/insights/ has been introduced to monitor these four practices. Although the Circle CI documentation references DORA metrics, there is no clear indication of support for monitoring them. The documentation also cites several third-party tools that can be utilized to monitor DevOps practices, such as Datadog, Honeycomb, Sumo Logic, Lightstep, New Relic, Splunk, Graylog, ELK stack, and metrics-capturing tools like Prometheus, Nagios, and Zabbix.
@@
AppVeyour offers a range of integration with various tools, but these integration mainly cover basic build metrics such as “Build Status” and “Build Duration”. Azure monitoring encompasses a complex set of services; however, its focus primarily lies in cloud infrastructure monitoring.
@@
Many monitoring tools, such as Graylog, Zabbix, Prometheus, and the ELK stack, are generic monitoring tools that can be used for CI monitoring but do not implement monitoring of any specific practices. As the ELK stack documentation describes: “ ELK stack obtain data reliably and securely from any source, in any format; then perform searches, analyzes and visualizations.”
@@
In conclusion, our findings indicate that Gitlab stands out as the most comprehensive CI service for monitoring CI practices. While DataDog also offers an extensive monitoring suite. Despite their individual strengths, the combination of these tools, still falls short of fully supporting CI practices monitoring. Moreover, they monitor some identical practices, such as Coverage, DORA metrics, Build Duration and Build Health. The integration of both tools introduces challenges as it unveils redundant monitoring practices, resulting in duplicated features and increased financial overhead for companies. Figure 16 provides an overview of CI tool areas. As we can see, the majority of tools concentrate on monitoring the CI pipeline practices, while some focus on aspects related to source code, such as test writing and coverage. Few tools address both of these practices groups. However, there is a notable absence of tools that specifically target monitoring practices related to the developer, such as commit frequency, commit size and comments per change. We only found an indication of monitoring commit frequency (commit activity) by Gitlab.
Figure 16: CI Tools Areas
Moreover, we realize that we can divide the most of practices $3233\frac${32}{33}divide start_ARG 32 end_ARG start_ARG 33 end_ARG in two groups regarding their purpose: Productivity (Speed) and Resilience (Quality). Figure 17 outlines this division.
Figure 17: DevOps Practices purpose
In Table 9, we establish a relationship between the DevOps practices documented in the CI services’ official documentation and the seven CI practices defined in the literature by Duvall et al Duvall2007 . Most of the DevOps practices can be linked to one or more CI practices defined by Duvall et al Duvall2007 . For instance, the practices “deploy frenquency” and “cycle time” are associated with the CI practice defined by Duvall et al as “Don’t commit broken code”. Don’t commit broken code helps to increase the “deploy frenquency” and “cycle time”. Another example is when you ’Write automated developer tests’, which contributes to increasing the “uptime” of the application and reduce the “change failure rate”.
@@
Table 9: Relation between Duvall CI practices Duvall2007 and DevOps practices
Duvall CI practice DevOps practice
@@
In Figure 18, we present an overview divided by DevOps area, showcasing the state-of-the-art monitoring of DevOps practices compiled from the official documentation of the CI services and third-party monitoring tools.
Figure 18: DevOps practices monitoring overview

### v2

@@ 16 Vulnerability scan of the resulting docker image 1
Our findings corroborate the results from RQ1. In RQ1, we found that $237285\frac${237}{285}divide start_ARG 237 end_ARG start_ARG 285 end_ARG (or 83%) of the monitoring opportunities suggested the monitoring of “Coverage,” of which $205237\frac${205}{237}divide start_ARG 205 end_ARG start_ARG 237 end_ARG (or 86%) were associated with the CodeCov tool. Sonarcloud was the third most used tool among survey respondents, which is probably why “Code Quality” and “Security” practices represent the second and third strongest evidence of monitoring in PR comments, with respectively 25 and 17 mentions. Code quality and security are typically monitored by tools such as Sonarcloud.
Evaluation of CI metrics
@@
When asked about the importance of monitoring the CI metrics presented in our survey (see Figure 3), most participants $2028\frac${20}{28}divide start_ARG 20 end_ARG start_ARG 28 end_ARG (or 71%), provided favourable responses. Developer 7 stated: “Yes, these metrics will provide valuable information in the status of our project over time.” Developer 15: “yes, it can make the development more efficient with high quality”, Developer 24: “Yes, especially build duration, build health, and time to fix broken builds”, Developer 20: “Absolutely, it would help show the health of the project, both for the active maintainers, but also for new people evaluating the project to determine if they want to use it..”
@@
Other participants found the metrics useful but not critical for their projects. Developer 9 mentioned: “In other projects these measurements are important, but in this Open Source project most of the contributors just work on the project when they have some spare time. The information is for sure useful, but not critical for us.” Some participants found only part of the metrics useful. Developer 25 mentioned: “depends on the practice, for example ‘commit frequency’ is not important at all, but fixing a broken build is.” Only 2 developers found the information about CI metrics evolution not useful at all. Developer 26 expressed: “Not really. I don’t think they are important because I know what needs to get done and time is the biggest limit.”, while Developer 21 said an emphatic “No.”
@@
We presume that Developer 21 is a relatively inexperienced developer. He stated that he has 5-10 years of experience with software development but is not involved in the maintenance of the CI infrastructure of the project, and he does not know which tools are involved in the project’s CI. He classified his project as having a “Base” level of CI maturity. However, because he was not involved in the project’s CI configurations and is not familiar with the CI tools used, his answers may not be entirely reliable.
@@
Developer 26 has 10-15 years of experience with software development and classified his project as having an “Intermediate” CI maturity level. His project is an embedded JavaScript library designed to include a special kind of graphics in html pages and provide them with extra functionality. Due to the nature of the project (e.g., it is not a high-scale critical project), following all the CI practices may not be deemed necessary, what justifies his answer.
@@
Almost half of respondents $124\frac{11}${24}divide start_ARG 11 end_ARG start_ARG 24 end_ARG (or 46% ) would like to improve their CI metrics. Developer 8 stated: “Absolutely! It saves time in the long run and increases confidence when refactoring is applied.”, Developer 28: “Yes, especially build time, fast CI = fast tests and deploy”, Developer 19: “Yes. Noticing problems early generally requires less engineering time and creates less distractions compared to noticing quality problems later.”, Developer 11: “Build time is something we should definitely worth dedicating time to. Coverage seems okay for us. Comments per change is not that much of an important aspect for us. We require at least two reviewers to approve and usually at our scale, we are positive that the reviewers usually do a good job.”, Developer 2: “Yes. Tracking the evolution of practices over time helps a project identify areas for improvement, evaluate changes, ensure consistency, and demonstrate progress to stakeholders.” and Developer 4: “Yes, as it affects the quality of our software product.”
@@
While some participants believe that it is important to invest time in improving the CI metrics presented in our survey, others $524\frac${5}{24}divide start_ARG 5 end_ARG start_ARG 24 end_ARG ( or 21% ) do not consider it a priority. For example, Developer 22 stated: “Sure, if someone has time”, while Developer 9 explained: “In this open source project, it is not critical because we have people who tend to fix issues promptly. However, in teams creating software for a company, it is much more important to keep the team productive and avoid being hindered by failures in CI.”
@@
Some participants $524\frac${5}{24}divide start_ARG 5 end_ARG start_ARG 24 end_ARG ( or 21% ) do not believe that improving these CI metrics are important. Developer 16 explained: “For this project, it doesn’t make much sense because the current approach seems to be working well”, while Developer 26 stated: “I don’t see how these metrics are connected to important goals like reducing errors”
@@
When asked if they find regular updates on the CI metrics useful, 68% of the developers $1928\frac${19}{28}divide start_ARG 19 end_ARG start_ARG 28 end_ARG answered “Yes”. The responses indicate that the majority of developers consider the studied CI metrics important and would monitor them if provided with the opportunity.
@@
In Figure 11, we code the benefits listed by developers when we asked about the importance of monitoring the CI metrics presented in our survey. They were grouped in three main themes: (i) “benefits for the project”, (ii) “benefits for the product” and (iii) “benefits for the process”. We recall that the coding process was simpler and differed from those used in the Document Analysis of RQ1 and RQ3, as it was not reviewed by other authors. The objective was to summarize the benefits listed by developers, in a single and simplified view.
Figure 11: Benefits of monitoring CI metrics
@@
Our findings related to this theme also corroborates our results from RQ1, where we primarily identified opportunities to monitor “build health” and “time to fix a broken build.” These two metrics were considered important by developers. For example, Developer 24 stated: “Yes, especially build duration, build health, and time to fix broken builds, while Developer 20 mentioned: “Absolutely. It would help show the health of the project, both for the active maintainers and also for new people evaluating the project to determine if they want to use it.”
@@
Although most developers feel that CI impacts productivity and quality of projects, few developers $218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) found the statistical correlations between CI metrics and bug-related issues or number of closed PRs useful (see Figure 4). Developer 16: “Hard to say but it makes sense that more commits lead to more closed pull requests” and Developer 13: “Generally one commit = one pull request, so I don’t think this particular metric means anything for this project.” Developer 24: “No, I did not find these 3 correlations interesting or meaningful. Not sure it will help somehow.”
@@
Only 2 developers perceived advantages in the correlations between CI metrics and quality attributes presented in the survey: Developer 28: “Yes, diagnostic reasons” and Developer 20: “Yes, it shows that the CI pipeline is helpful.”
@@
In this last part of the survey, we asked participants whether there were any CI metrics we missed or whether our participants preferred other types of visualizations for the metrics presented.
@@
$218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) developers expressed that the metrics and visualizations presented were adequate. Developer 2: ‘Line charts can show trends in practice performance over time, while bar charts can compare the performance of different practices. Both are useful for a report on CI practices.” and Developer 27: “graphs were perfect.” $318\frac${3}{18}divide start_ARG 3 end_ARG start_ARG 18 end_ARG (or 16% ) did not miss any information, they answered a simple: “No.” $218\frac${2}{18}divide start_ARG 2 end_ARG start_ARG 18 end_ARG (or 11% ) developers are not sure which metrics they would like to see monitored. Developer 10: “Not sure” and Developer 22: “I’m not sure.” $118\frac${1}{18}divide start_ARG 1 end_ARG start_ARG 18 end_ARG (or 5% ) Developer provided an unhelpful answer and has been discarded: “CI environments are sometimes really annoying to debug.” $1018\frac${10}{18}divide start_ARG 10 end_ARG start_ARG 18 end_ARG (or 55% ) suggested improvements to the information contained in the survey. We describe some of the suggestions in the next few paragraphs.
@@
Developer 15, while liked the visualization of the metrics in the survey: “The pictures above including these 7 charts are interesting to me.”, suggested including security metrics like image scan, code dependency scan, and configuration scan. Although important, these security metrics are outside the scope of our survey as they do not directly relate to CI but rather to DevOps. Developer 28 suggested including “deploy time” as an interesting metrics for CD in future research.
@@
Participants also discussed other matters related to PR, Issue, and Commit data. Developer 11 stated, “PR and issue related data seems to be missing. It should be public, so you should be able to gather that from our GitHub project. We use tags to sort them.” Developer 13 suggested, “Perhaps commit size over time too? Tracking keywords used in commit messages?.” Developer 12 also like how the information of seven CI metrics were presented in the survey: “I think the above information is enough. For our current usage, having these simple graphs available for a daily/weekly tracking of the project status would be sufficient to assign someone to fixing the problem. We do this in a more informal way currently”, but recommended investigating “statistics on who reviews and who commits, if there’s a change in the balance (less committers or less reviewers) as important. The aim is to ensure knowledge sharing, and that one developer does not ’take over the project’ while breaking the information and knowledge sharing that is in place.” While PR and Issue information were used to compute correlations regarding CI practices in our study, they are not CI metrics and were not the focus of this paper. Both “Commit size” and “Ratio between committers and reviewers,” as well the consequences of an unbalanced ratio, are interesting aspects to investigate in future research.
@@
For a broader view of the monitoring needs, in Figure 12 we show the themes that emerged from the monitoring needs suggested by survey participants. As same way of Figure 11, we recall that the coding process was simpler and differed from those used in the Document Analysis of RQ1 and RQ3. The objective was to summarize the monitoring needs in a single and simplified view.
Figure 12: Needs related to monitoring CI metrics
@@
Figure 13 illustrates the themes that emerged from the official documentation of the seven most popular CI services GolzadehCIServer2022 . We emphasize that the results described below are derived from the thematic analysis conducted on the official documentation of the CI services included in this study and are limited to the metrics present in the examined documents.
@@
In this analysis we find other metrics mentioned outside the scope of CI, they were classified under the umbrella of “DevOps metrics”. For example, the DORA metrics DoraMetricsLeanIX2023 have been widely mentioned for the CI services. Gitlab, GHActions, and Circle CI cite these metrics. However, only the Gitlab documentation declare that there is an implementation of a dashboard to monitor these metrics. In May 22, 2023141414https://about.gitlab.com/releases/2023/05/22/gitlab-16-0-released/, Gitlab has added the “Value Stream Dashboard” which included the DORA and Vulnerabilities metrics, but is still accessible only from Premium version [^15] https://stackoverflow.com/questions/76382478/permission-denied-for-gitlab-dora-metrics-through-api. Gitlab was the service with the highest number of citations ($5161\frac{55}${161}divide start_ARG 55 end_ARG start_ARG 161 end_ARG) (or 34% ) for themes that refer to DevOps metrics monitoring, 1.53 times more than the second place CicleCI with ($3161\frac{33}${161}divide start_ARG 33 end_ARG start_ARG 161 end_ARG) (or 20% ). This indicates that Gitlab is the most advanced CI service in terms of supporting the monitoring of DevOps metrics, without the need to use third-party tools.
In the GHActions documentation, we did not find indications of native monitoring tools specifically designed for CI metrics. The existing monitoring tools appear to concentrate on optimizing the CI/CD pipeline by monitoring aspects such as CPU and memory usage within the pipeline. However, GHActions allows the creation of various plugins [^16] https://github.com/marketplace?type=apps&category=continuous-integration that can complement these monitoring tasks.
Circle CI focuses on monitoring four metrics: 1) “Build Health”, 2) “Mean Time to Recovery”, 3) “Build Duration”, and 4) “Build Activity”. In January 2022, a centralized dashboard [^17] https://circleci.com/docs/insights/ has been introduced to monitor these four metrics. Although the Circle CI documentation references DORA metrics, there is no clear indication of support for monitoring them. The documentation also cites several third-party tools that can be utilized to monitor DevOps metrics, such as Datadog, Honeycomb, Sumo Logic, Lightstep, New Relic, Splunk, Graylog, ELK stack, and metrics-capturing tools like Prometheus, Nagios, and Zabbix.
@@
Appveyor offers a range of integration with various tools, but these integration mainly cover basic build metrics such as “Build Status” and “Build Duration”. Azure monitoring encompasses a complex set of services; however, its focus primarily lies in cloud infrastructure monitoring.
@@
Many monitoring tools, such as Graylog, Zabbix, Prometheus, and the ELK stack, are generic monitoring tools that can be used for CI monitoring but do not implement monitoring of any specific metrics. As the ELK stack documentation describes: “ ELK stack obtain data reliably and securely from any source, in any format; then perform searches, analyzes and visualizations.”
@@
In conclusion, our findings indicate that Gitlab stands out as the most comprehensive CI service for monitoring CI practices. While DataDog also offers an extensive monitoring suite. Despite their individual strengths, the combination of these tools, still falls short of fully supporting CI practices monitoring. Moreover, they monitor some identical metrics, such as Coverage, DORA metrics, Build Duration and Build Health. The integration of both tools introduces challenges as it unveils redundant monitoring metrics, resulting in duplicated features and increased financial overhead for companies. Moreover, we realize that we can divide the most of metrics $3233\frac${32}{33}divide start_ARG 32 end_ARG start_ARG 33 end_ARG in two groups regarding their purpose: Productivity (Speed) and Resilience (Quality). Figure 15 outlines this division.
Figure 15: DevOps Metrics purpose
In Table 10, we establish a relationship between the DevOps metrics documented in the CI services’ official documentation and the seven CI practices defined in the literature by Duvall et al Duvall2007 . Most of the DevOps metrics can be linked to one or more CI practices defined by Duvall et al Duvall2007 . For instance, the metrics “deploy frequency” and “cycle time” are associated with the CI practice defined by Duvall et al as “Don’t commit broken code”. Don’t commit broken code helps to increase the “deploy frequency” and “cycle time”. Another example is when you ’Write automated developer tests’, which contributes to increasing the “uptime” of the application and reduce the “change failure rate”.
@@
Table 10: Relation between Duvall CI practices Duvall2007 and DevOps metrics
Duvall CI practice DevOps metrics
@@
In Figure 16, we present an overview divided by DevOps area, showcasing the state-of-the-art monitoring of DevOps metrics compiled from the official documentation of the CI services and third-party monitoring tools.
Figure 16: DevOps metrics monitoring overview